OpenStudy (anonymous):

integral of 2(tan x)(sec x)^2 dx = (tan x) ^2 = (sec x)^2? is this so?

OpenStudy (anonymous):

Secant x and Tangent x are not the same thing, so no.

OpenStudy (raden):

int 2 (tanx) (secx)^2 dx = 2 int (tanx secx) secx dx use int u-sub let u = secx du = secx tanx dx see ur integral becomes 2 int u du = ... does that helps ?

OpenStudy (anonymous):

\[\int\limits_{}^{}2 \tan x \sec ^{2}x dx \] if I put u = tan x, du = \(\sec ^{2} x dx\), so it will be \(\int\limits_{}^{}2u du= \frac{ 2u ^{2} }{ 2 }=u ^{2} = \tan ^{2}x\) but if I use u=sec x, du = sec x tan x dx, so I get \(\int\limits_{}^{}2udu = \frac{ 2u ^{2} }{ 2 } = \sec^{2}x\) Did I do something wrong?

OpenStudy (raden):

no, u are right the result is same, just different for the constant

OpenStudy (anonymous):

oh I see, that's pretty interesting.

OpenStudy (raden):

yeah, if we derive it then will back to the original function :)